Shannon purchased a $4,500 bond that was paying a coupon rate of 6.70% compounded semi-annually and had 4 more years to mature. The yield at the time of purchase was 5.30% compounded semi-annually.

a. How much did Shannon pay for the bond?

What was the amount of premium or discount on the bond?  

Premium or Discount? and amount was _
Answer #1

a.

Bond Par Value, FV = $4,500

Semi Annual Coupon, PMT = (0.067/2)(4,500) = $150.75

Time Period, N = 8 semi-annual Periods

YTM = 0.053

Calculating Bond Price,

Using TVM Calculation,

PV = [FV = 4,500, PMT = 150.75, N = 8, I = 0.053/2]

PV = $4,724.42

So,

Bond Price = $4,724.42

As Bond Price is greater than par value, Bond is trading at premium to par value,

Premium = 4,724.42 - 4,500

Premium = $224.42